Solve
(5/z)=(7/z)-(1/5)
I don't understand how the solution can end = 5/z

Answer by stanbon(75887)

Solve
(5/z)=(7/z)-(1/5)
I don't understand how the solution can end = 5/z
----
Multiply thru by 5z to get:
25 = 35 - z
z = 10
=============
Cheers,
stan H.